# Evaluating Concurrent Robustness of Language Models Across Diverse Challenge Sets

### v2

@@ 1 Introduction
Our proposed methodology is adept at handling both (a) transformer-based pre-trained language models (PLMs) such as BERT (Devlin et al., 2018) and RoBERTa (Liu et al., 2019c) , which are amenable to direct fine-tuning, and (b) large generative language models such as gpt-3.5-turbo(GPT-3.5) (Brown et al., 2020), GPT-4, and LLaMA, LLaMA-2 (Touvron et al., 2023), Flan-T5(Chung et al., 2022; Kanakarajan and Sankarasubbu, 2023) etc, which can’t be fine-tuned freely. For these generative models, we leverage the few-shot Chain of Thought (Wei et al., 2023) as an alternative to traditional fine-tuning. This methodology circumvents the computational intricacies inherent in the fine-tuning of LLMs. It proficiently manages the tuning of a multitude of model parameters using a limited constrained set of training samples. To the best of our knowledge, we are the first to study Inoculation with LLM, prior studies Liu et al. (2019c); Wang et al. (2021a); Liu et al. (2019b) have been limited to traditional BERT style models. Within our framework, we investigate three distinct multi-set fine-tuning methods, each designed to enhance model robustness across diverse perturbation sets. Our study makes the following contributions:
@@ 2 Proposed Methodology
Terminology. Given a pre-trained language model (PLM) denoted as M, fine-tuned on the original (unperturbed) training set O for a natural language processing (NLP) task T. Let {} represent input perturbations, where is the number of distinct perturbations available. For each perturbation , let O be a subset Sj of the original training set O, where . Let represent an input perturbation applied only to O, producing the perturbation/challenge set . This results in perturbation sets , where perturbation is applied to subset , respectively. We use Pj as shorthand for the final perturbation set . We evaluate the performance of model M on held-out perturbation set samples Qj for . Each Qj serves as the test set specifically tailored for perturbation .
@@ 2.2 Single Model Multi Set Inoculation
To alleviate the complexity of managing multiple robust models, we propose cultivating a universal robust model(RM) that remains immune to various perturbations in input data. We put forth three distinct fine-tuning strategies for the same: Sequential (Seq): The model is fine-tuned using k samples from each challenge set Pj sequentially (specified by fixed Order), resulting into a final robust model RM. Mixed-Training (Mix): In this strategy, a composite dataset, termed PM, is fashioned by randomly selecting k samples from all challenge sets , {Pj}. Subsequently, the model M is fine-tuned using the aggregated PM. In our implementation, we adopt a uniform, random sampling approach. Dynamic Mix-Training (DynMix): This approach mirrors mixed-training but introduces variability in sample sizes across different challenge sets, denoted as k1, k2, and so on. Additionally, the sampling method can be unique (e.g. uniform or weighted) for each perturbation challenge set.
@@
(a) Zero-shot (OP). We create a prompt template consisting of only the description of the task, without any exemplars or reasoning chains. (b) Few-shot with CoT (OP). Here, we consider NLI task description along with few shot exemplars taken from the original set O their reasoning chains a.k.a. COT.
Single Exemplars Multiple Prompts (SEMP): For each perturbation type, denoted as , we construct a prompt that combines the task description, respective perturbation description, and exemplars from O and Pj. The exemplars are accompanied by corresponding labels and a reasoning chain (CoT). This results in multiple prompts, each tailored to a specific perturbation . We call this approach single exemplars multiple prompts, similar to multi model single set in sec. 2.1.
Multiple Exemplars Single Prompt (MESP) : Here, we consider descriptions and exemplars of all perturbations () in a single prompt. We create a prompt by combining multiple exemplars corresponding to each perturbation , sampled from Pj, similar to single model multi set in section 2.2. Here, the prompt contains the task description, a description of all perturbations, and exemplars from the original set O and each of the challenge sets ( Pj). Given token length constraints, tradeoff between the detail of perturbation descriptions and the number of perturbation exemplars results in two variants: (a) Mixed-Prompting-Instructional (MESP) : In this prompt, the perturbation description is emphasized while reducing the number of exemplars. Mixed-Prompting-Exemplar (MESP): Here more perturbation exemplars are sampled and each perturbation’s description is shortened.
@@ 3 Case Study on Tabular Inference
Original Dataset (O). We utilize the tabular-NLI dataset, InfoTabS (Gupta et al., 2020), along with its adversarial perturbations, as detailed in Shankarampeta et al., 2022. The InfoTabS dataset features a wide range of table domains, categories, and keys, covering various entity types and forms. It includes three test splits: (original test set), (adversarial set), and (zero-shot or out-of-domain set).
Perturbed Challenge Datasets (P, Q). Our dataset incorporates perturbations from Shankarampeta et al., 2022, enhanced using tools such as TextAttack (Morris et al., 2020a) and NLP Checklist (Ribeiro et al., 2020), alongside manual adjustments. Each perturbation specifically targets the hypothesis of an input sample. For every perturbation type, we create challenge sets of up to 1,500 samples. Only those samples that are pertinent post-perturbation are selected. When the number of such samples exceeds 1500, we narrow down to the most diverse 1500 samples using Fixed-Size Determinantal Point Processes (-DPPs) (Kulesza and Taskar, 2011). Perturbations used for Tabular-NLI tasks are Character-level perturbation (char, C), Negation-type perturbation (neg, N), Numeric perturbation (num, M), Location perturbation (loc, L) and Paraphrasing perturbation (stan, S) (refer Figure 1).
Train/Test. (a.) BERT Based Models (PLM) : For any perturbation type, we represent Qj consisting of 1000 examples for testing and Pj consisting of 500 examples for fine-tuning. We define the union of all challenge test sets as and the training set as . (b.) Large Language Models (LLM) : As LLMs inference is costly we limit our evaluations to 300 random samples from , where contains original premise and perturbed hypothesis using perturbation . contains the original premise along with the corresponding unperturbed hypothesis as pairs. We evaluate performance on both and to access if the LLM model forgets the original input distribution after fine-tuning on perturbation sets.
Table Representation. In line with Neeraja et al., 2021, we employed alignment techniques in Yadav et al., 2020 to eliminate distracting rows (DRR). We selected the top-8 rows for table representation as a premise (DRR@8), enhancing accuracy through evidence-based grounding of relevant information for hypothesis labeling.
@@ 3.1 Fine-tuning BERT Based Model
Sequential (Seq): We perform sequential fine-tuning of RoBERTa across various challenge sets. The training order (Order) for fine-tuning is based on average baseline model performance across challenge sets. Our sequencing strategy aims to minimize the potential for catastrophic forgetting (Kirkpatrick et al., 2017; Goodfellow et al., 2013) induced by subsequent fine-tuning on challenge sets.
Mixed-Training (Mix): Here, the RoBERTa is fine-tuned samples obtained by mixing k instances drawn from each of the challenge sets . Here, k is an hyper-parameters, set equal to examples, as discussed in section 3.1.
@@ 3.2 LLM Prompting
We used GPT-3.5 with low temperature of , LLaMA-2 after quantization using QLoRA (Dettmers et al., 2023), and Flan-T5 series. We develop methodologies for LLMs that rely solely on prompting and exclude fine-tuning (except for GPT-3.5 where we also report fine-tuning results). The LLM prompt design for our experiments, is detailed in Table 1, comprises five sections, with demonstration section being optional.
@@
| Answering | (Restriction for Answering) Answer with an explanation in the following format, restricting the answer to only one of the following: "yes" or "no" or "it is not possible to tell" + Answering Format |
@@ 4.1 Results: Bert Style Models (PLM)
Terminology. To define the sequence we consider (a.) Column Wise Average. This configuration assesses the aggregate impact of fine-tuning across all perturbations on each individual perturbation., (b.) Row Wise Average. This configuration evaluates the aggregate impact of fine-tuning on an individual perturbation against all other perturbations. For more details on the metrics refer to Appendix A.3.
@@
Ablation Experiments. (a) Fine tuning on subset of Perturbation. Above Mix and DynMix requires access to all perturbation during fine-tuning, which increasing dataset and computation cost. To access whether robust models can be obtained via fine-tuning on a subset of perturbation sets, we ran experiments using subset of perturbations. The results are shown in Appendix A.1. Our results show that although there are performance improvements while fine-tuning on subsets of perturbation. Nevertheless, the optimal subset of available perturbations for the task remains elusive and cannot be found empirically. (b) Results on Out of Distribution Perturbations. Assessing the model’s performance against unseen perturbations is vital for robustness. Such evaluation reveals the model’s ability to adapt to new and unexpected changes. We created approximately 100 samples (with nearly equal numbers of E, C, N labels) of a new Word-Swap perturbation type. The results are shown in Appendix A.1. We observe fine-tuning with more samples using the Mix strategy enhances model robustness against unseen perturbations, further validating our approach.
@@ 4.2 Results: Large Language Models (LLMs)
Analysis. • On the Original Zero-Shot Prompts we observe that, (a.) Comparing the results of challenge datasets Qj and their unperturbed version sets Q’j reveals that LLMs similar to PLMs are also sensitive to input data perturbations. (b.) However, the Flan-T5 series, specifically XL and XXL, performs significantly better than other LLMs as it’s fine-tuned specifically for the NLI task (Chung et al., 2022). Even the drop in performance due to data perturbation is relatively less. (c.) The poor performance of relatively smaller LLMs, such as LLaMA-2-13b, demonstrates the ineffectiveness of such models in responding to an instruction prompt. (d.) One reason for performance on original numerical set (Q’M), is due to model inability to handle mathematical reasoning (Wallace et al., 2019; Min et al., 2021; Hendrycks et al., 2021; Imani et al., 2023). Additionally, we find that all models enhanced with CoT (Table 4) outperform those using Zero Shot original prompts. This suggests that simply adding exemplars can enhance a model’s resilience to perturbations.
Single Exemplars Multiple Prompts (SEMP): Table 5(a) presents results for GPT-3.5, with diagonal elements as an analogue to single set inoculation. LLaMA-2 results are in Table 5(b).
@@
| Strategy | char | neg | num | loc | stan | ||
| LLaMA-2 | Base | 39.67 | 39.33 | 45.67 | 56.67 | 44.67 | 45.20 |
| MESP | 58.33 | 51.33 | 59.67 | 55.00 | 62.33 | 62.53 | |
| MESP | 60.97 | 56.67 | 59.33 | 60.67 | 62.00 | 65.67 | |
| GPT-3.5 | Base | 51.00 | 53.00 | 62.66 | 61.00 | 60.30 | 69.45 |
| MESP | 59.60 | 65.30 | 62.00 | 59.00 | 59.60 | 68.76 | |
| MESP | 70.00 | 63.60 | 58.30 | 60.00 | 66.80 | 72.52 |
Multiple Exemplars Single Prompts (MESP): The results for MPI and MPE are in Table 6.
@@
Fine-tuning on LLMs. While our paper primarily focuses on in-context learning for LLMs, we also examine the effects of fine-tuning LLMs on perturbation sets, results shown in Table 7. We can see that for Mixtral and GPT-3.5 the fine tuning with the perturbation set using the mix training approach increases the models’ performance. Whereas for the Flan-T5-L model the fine tuning does not improve the model’s performance.
| Model | char | neg | num | loc | stan | avg. | |
| BASE | Flan-t5-L | 63.00 | 70.00 | 63.00 | 65.00 | 69.30 | 66.06 |
| Mistral | 44.01 | 39.66 | 23.67 | 41.34 | 39.67 | 37.67 | |
| GPT-3.5 | 51.00 | 53.00 | 62.66 | 61.00 | 60.30 | 57.59 | |
| FT | Flan-t5-L | 51.71 | 27.77 | 45.31 | 42.71 | 58.70 | 45.24 |
| Mistral | 43.67 | 39.00 | 61.67 | 45.67 | 44.34 | 46.87 | |
| GPT-3.5 | 71.67 | 78.33 | 66.67 | 67.33 | 71.00 | 71.00 |
@@ 5 Related Works
Model Robustness Issues. Deep learning models in vision and language domains have exhibited sensitivity to adversarial examples and input distribution shifts, as highlighted in prior studies Mahmood et al. (2021); Elsayed et al. (2018); Chang et al. (2021); Ren et al. (2019); McCoy et al. (2019); Wang et al. (2021a); Gupta et al. (2023); Zheng and Saparov (2023); Zhu et al. (2023). The search for model robustness in language processing has led to work on contrast sets (Li et al., 2020a), Checklist (Ribeiro et al., 2020), and attack algorithms (Li et al., 2020b, 2018). Ensuring model robustness is crucial (Wang et al., 2022, 2020), as minor input changes can significantly impact performance due to model complexity and distribution overfitting (Glockner et al., 2018; Rice et al., 2020; Zhu and Rao, 2023; Moradi and Samwald, 2021). Recently, Zhu et al. (2023) introduce adversarial prompts to analyse model robustness to perturbatin in prompts. Our paper focuses on analyzing model performance with clean prompt across several perturbations/attacks on input samples simultaneously.
Improving Model Robustness. Utilizing adversarial examples during training provides a degree of mitigation (Tong et al., 2022; Liu et al., 2019a; Yuan et al., 2023; Kotha et al., 2023; Liu et al., 2023), it falls short of a comprehensive solution for achieving widespread robustness, as it deals only with one facet, i.e., single-set inoculation. Our proposed framework is adept at evaluating model robustness across multiple challenge sets. Our research complements and extends the work on robustness explored in Liu et al. (2023); Lu (2022); Zheng and Saparov (2023). While Liu et al. (2023) integrates consistency loss and data augmentation during training, our framework applies to models already in use or deployed. Similarly, Lu (2022) addresses dataset artifacts in natural language inference (NLI) with a multi-scale data augmentation method. In contrast, our work focuses on limited fine-tuning of pre-trained models and expands to additional dimensions of robustness. Meanwhile, Zheng and Saparov (2023) examines LLM robustness to perturbed inputs by increasing noisy exemplars. Our study offers a broader framework for assessing the robustness of both PLMs and LLMs, using fine-tuning, improving instruction quality, and enhancing exemplars in both diversity and quantity.

### v3

@@ 1 Introduction
Our proposed methodology is adept at handling both (a) transformer-based pre-trained language models (PLMs) such as BERT (Devlin et al., 2018) and RoBERTa (Liu et al., 2019c) , which are amenable to direct fine-tuning on end-user GPUs, and (b) large generative language models such as gpt-3.5-turbo (GPT-3.5) (Brown et al., 2020), GPT-4, and LLaMA, LLaMA-2 (Touvron et al., 2023), Flan-T5 (Chung et al., 2022; Kanakarajan and Sankarasubbu, 2023), which are costly and have limited access to re-training (and model weights). For these generative models, we leverage the few-shot Chain of Thought (Wei et al., 2023) as an alternative to traditional fine-tuning. This methodology circumvents the computational intricacies inherent in the fine-tuning of LLMs. It proficiently manages the tuning of a multitude of model parameters using a limited constrained set of training samples. Additionally, we also study Inoculation with LLM, prior studies Liu et al. (2019c); Wang et al. (2021a); Liu et al. (2019b) have been limited to traditional BERT style models. Within our framework, we investigate three distinct multi-set fine-tuning methods for PLMs and adapt them for LLMs via COT, each designed to assess and enhance model robustness across diverse perturbation sets. Our study makes the following contributions:
@@ 2 Proposed Methodology
Terminology.
Given a pre-trained language model (PLM) denoted by M, fine-tuned on the original (unperturbed) training set O for a natural language processing (NLP) task T.
where represent the sample-label pair in the dataset. Let represent input perturbations, where is the number of distinct perturbations available. For each perturbation , let O be a subset of the original training set O, where :
Perturbation is applied only to O, producing the perturbation/challenge set :
This results in perturbation sets where perturbation is applied to subset . We use Pj as shorthand for the final perturbation set . We evaluate the performance of model M on held-out perturbation set samples Qj for . Each Qj serves as the test set specifically tailored for perturbation .
@@ 2.2 Single Model Multi Set Inoculation
To alleviate the complexity of managing multiple robust models, we propose cultivating a universal robust model(RM) that remains immune to various perturbations in input data. We put forth three distinct fine-tuning strategies for the same:
Sequential (Seq): The model is fine-tuned using k samples from each challenge set Pj sequentially (specified by fixed Order), resulting into a final robust model RM.
Mixed-Training (Mix): In this strategy, a composite dataset, termed PM, is fashioned by randomly selecting k samples from all challenge sets , {Pj}. Subsequently, the model M is fine-tuned using the aggregated PM. In our implementation, we adopt a uniform, random sampling approach.
Dynamic Mix-Training (DynMix): This approach mirrors mixed-training but introduces variability in sample sizes across different challenge sets, denoted as k1, k2, and so on. Additionally, the sampling method can be unique (e.g. uniform or weighted) for each perturbation challenge set.
@@
(a) Zero-shot (OP). We create a prompt template consisting of only the description of the task, without any exemplars or reasoning chains.
(b) Few-shot with CoT (OP). Here, we consider NLI task description along with few shot exemplars taken from the original set O their reasoning chains a.k.a. CoT.
Single Exemplars Multiple Prompts (SEMP): For each perturbation type, denoted as , we construct a prompt that combines the task description, respective perturbation description, and exemplars from O and Pj. The exemplars are accompanied by corresponding labels and a reasoning chain (CoT). This results in multiple prompts, each tailored to a specific perturbation . We call this approach single exemplars multiple prompts, similar to multi model single set (refer sec. 2.1).
Multiple Exemplars Single Prompt (MESP) : Here, we consider descriptions and exemplars of all perturbations () in a single prompt. We create a prompt by combining multiple exemplars corresponding to each perturbation , sampled from Pj, similar to single model multi set in section 2.2. Here, the prompt contains the task description, a description of all perturbations, and exemplars from the original set O and each of the challenge sets ( Pj). Given token length constraints, a trade-off between the detail of perturbation descriptions and the number of perturbation exemplars results in two variants:
(a) Mixed Prompting Instructional (MESP): In this prompt, the perturbation description is emphasized while reducing the number of exemplars.
(b) Mixed Prompting Exemplar (MESP): Here more perturbation exemplars are sampled and each perturbation’s description is shortened.
@@ 3 Case Study on Tabular Inference
Original Dataset (O).
We utilize the tabular-NLI dataset, InfoTabS (Gupta et al., 2020), along with its adversarial perturbations, as detailed in Shankarampeta et al., 2022. The InfoTabS dataset features a wide range of table domains, categories, and keys, covering various entity types and forms. It includes three test splits: (original test set), (adversarial set), and (zero-shot or out-of-domain set).
Perturbed Challenge Datasets (P, Q).
Our dataset incorporates perturbations from Shankarampeta et al., 2022, enhanced using tools such as TextAttack (Morris et al., 2020a) and NLP Checklist (Ribeiro et al., 2020), alongside manual adjustments. Each perturbation specifically targets the hypothesis of an input sample. For every perturbation type, we create challenge sets of up to 1,500 samples. Only those samples that are pertinent post-perturbation are selected. When the number of such samples exceeds 1500, we narrow down to the most diverse 1500 samples using Fixed-Size Determinantal Point Processes (-DPPs) (Kulesza and Taskar, 2011). Perturbations used for Tabular-NLI tasks are Character-level perturbation (char, C), Negation-type perturbation (neg, N), Numeric perturbation (num, M), Location perturbation (loc, L) and Paraphrasing perturbation (stan, S) (refer Figure 1).
Train/Test.
(a.) BERT Based Models (PLM) : For any perturbation type, we represent Qj consisting of 1000 examples for testing and Pj consisting of 500 examples for fine-tuning. We define the union of all challenge test sets as and the training set as .
(b.) Large Language Models (LLM) : As LLMs inference is costly we limit our evaluations to 300 random samples from , where contains original premise and perturbed hypothesis using perturbation . contains the original premise along with the corresponding unperturbed hypothesis as pairs. We evaluate performance on both and to access if the LLM model forgets the original input distribution after fine-tuning on perturbation sets.
Table Representation. In line with Neeraja et al., 2021, we employed alignment techniques (Yadav et al., 2020) to eliminate distracting rows (DRR). We selected the top-8 rows for table representation as a premise (DRR@8), enhancing accuracy through evidence-based grounding of relevant information for hypothesis labeling.
@@ 3.1 Fine-tuning BERT Based Model
Sequential (Seq): We perform sequential fine-tuning of RoBERTa across various challenge sets. The training order (Order) for fine-tuning is based on average baseline model performance across challenge sets. Sequential fine-tuning often leads to catastrophic forgetting of previously learned perturbations (Kirkpatrick et al., 2017; Goodfellow et al., 2013). To mitigate this, we propose two alternative strategies designed to minimize this effect.
Mixed-Training (Mix): Here, the RoBERTa is fine-tuned samples obtained by mixing k instances drawn from each of the challenge sets . Here, k is an hyper-parameter, set equal to examples, as discussed in section 3.1.
@@ 3.2 LLM Prompting
We used GPT-3.5 with low temperature of , LLaMA-2 after quantization using QLoRA (Dettmers et al., 2023), Mistral Jiang et al. (2023) and Flan-T5 series Chung et al. (2024).We develop methodologies for LLMs that rely solely on prompting and exclude fine-tuning (except for GPT-3.5 where we also report fine-tuning results). The LLM prompt design for our experiments, is detailed in Table 1, comprises five sections, with demonstration section being optional.
@@
| Answering | (Restriction for Answering) Answer with an explanation in the following format, restricting the answer to only one of the following: "yes" or "no" or "it is not possible to tell" + <Answering Format> |
@@
Complete prompt examples for each case can be found in Appendix A.3.
@@ 4.1 Results: Bert Style Models (PLM)
Terminology. To define the sequence we consider
(a.) Column Wise Average. This configuration assesses the aggregate impact of fine-tuning across all perturbations on each individual perturbation.
(b.) Row Wise Average. This configuration evaluates the aggregate impact of fine-tuning on an individual perturbation against all other perturbations. For more details on the metrics refer to Appendix A.3.
@@
Ablation Experiments.
(a) Fine tuning on a subset of Perturbations. Above Mix and DynMix require access to all perturbations during fine-tuning, which increases dataset and computation costs. To assess whether robust models can be obtained via fine-tuning on a subset of perturbation sets, we ran experiments using a subset of perturbations. The results are shown in Appendix A.1. Our results show that although there are performance improvements while fine-tuning on subsets of perturbation. Nevertheless, the optimal subset of available perturbations for the task remains elusive and cannot be found empirically.
(b) Results on Out of Distribution Perturbations. Assessing the model’s performance against unseen perturbations is vital for robustness. Such evaluation reveals the model’s ability to adapt to new and unexpected changes. We created approximately 100 samples (with nearly equal numbers of E, C, N labels) of a new Word-Swap perturbation type. The results are shown in Appendix A.1. We observe fine-tuning with more samples using the Mix strategy enhances model robustness against unseen perturbations, further validating our approach.
@@ 4.2 Results: Large Language Models (LLMs)
Analysis. • On the Original Zero-Shot Prompts we observe that, (a.) Comparing the results of challenge datasets Qj and their unperturbed version sets Q’j reveals that LLMs similar to PLMs are also sensitive to input data perturbations. (b.) However, the Flan-T5 series, specifically XL and XXL, performs significantly better than other LLMs as it’s fine-tuned specifically for the NLI task (Chung et al., 2022). Even the drop in performance due to data perturbation is relatively less. (c.) The poor performance of relatively smaller LLMs, such as LLaMA-2-13b, demonstrates the ineffectiveness of such models in responding to an instruction prompt. (d.) One reason for lower performance on original numerical set (Q’M), is due to model inability to handle mathematical reasoning (Wallace et al., 2019; Min et al., 2021; Hendrycks et al., 2021; Imani et al., 2023). Additionally, we find that all models enhanced with CoT (Table 4) outperform those using Zero Shot original prompts. This suggests that simply adding exemplars can enhance a model’s resilience to perturbations.
Single Exemplars Multiple Prompts (SEMP): Table 5(a) presents results for GPT-3.5, with diagonal elements as an analog to single set inoculation. LLaMA-2 results are in Table 5(b).
@@
Multiple Exemplars Single Prompts (MESP): The results for MPI and MPE are in Figure 4.
@@
Fine-tuning on LLMs. While our work primarily focuses on in-context learning for LLMs, we also examine the effects of fine-tuning LLMs on perturbation sets, results shown in Figure 5. We can see that for Mistral and GPT-3.5 the fine tuning with the perturbation set using the mix training approach increases the models’ performance. Whereas for the Flan-T5-L model the fine tuning does not improve the model’s performance.
@@ 5 Related Works
Model Robustness Issues.
Deep learning models in vision and language domains have exhibited sensitivity to adversarial examples and input distribution shifts, as highlighted in prior studies Mahmood et al. (2021); Elsayed et al. (2018); Chang et al. (2021); Ren et al. (2019); McCoy et al. (2019); Wang et al. (2021a); Gupta et al. (2023); Zheng and Saparov (2023); Zhu et al. (2023). The search for model robustness in language processing has led to work on contrast sets (Li et al., 2020a), Checklist (Ribeiro et al., 2020), and attack algorithms (Li et al., 2020b, 2018). Ensuring model robustness is crucial (Wang et al., 2022, 2020), as minor input changes can significantly impact performance due to model complexity and distribution overfitting (Glockner et al., 2018; Rice et al., 2020; Zhu and Rao, 2023; Moradi and Samwald, 2021). Recently, Zhu et al. (2023) introduce adversarial prompts to analyse model robustness to perturbation in prompts. Our work focuses on analyzing model performance with clean prompts across several perturbations/attacks on input samples simultaneously.
Improving Model Robustness.
Utilizing adversarial examples during training provides a degree of mitigation to input sensitivity of a deep learning model (Tong et al., 2022; Liu et al., 2019a; Yuan et al., 2023; Kotha et al., 2023; Liu et al., 2023), however, it falls short of a comprehensive solution for achieving widespread robustness, as it deals only with one facet, i.e., single-set inoculation. Our proposed framework is adept at evaluating model robustness across multiple challenge sets. Our research complements and extends the work on robustness explored in Liu et al. (2023); Lu (2022); Zheng and Saparov (2023). While Liu et al., 2023 integrates consistency loss and data augmentation during training, our framework applies to models already in use or deployed. Similarly, Lu, 2022 addresses dataset artifacts in natural language inference (NLI) with a multi-scale data augmentation method. In contrast, our work focuses on limited fine-tuning of pre-trained models and expands to additional dimensions of robustness. Meanwhile, Zheng and Saparov, 2023 examines LLM robustness to perturbed inputs by increasing noisy exemplars. Our study offers a broader framework for assessing the robustness of both PLMs and LLMs, using fine-tuning, improving instruction quality, and enhancing exemplars in both diversity and quantity.